### v1

# Parity-Aware Byte-Pair Encoding: Improving Cross-lingual Fairness in Tokenization

Abstract
Tokenization is the first—and often least scrutinized—step of most NLP pipelines. Standard algorithms for learning tokenizers rely on frequency-based objectives, which favor languages dominant in the training data and consequently leave lower-resource languages with tokenizations that are disproportionately longer, morphologically implausible, or even riddled with <UNK> placeholders. This phenomenon ultimately amplifies computational and financial inequalities between users from different language backgrounds. To remedy this, we introduce Parity-aware Byte Pair Encoding (BPE), a variant of the widely-used BPE algorithm. At every merge step, Parity-aware BPE maximizes the compression gain of the currently worst-compressed language, trading a small amount of global compression for cross-lingual parity. We find empirically that Parity-aware BPE leads to more equitable token counts across languages, with negligible impact on global compression rate and no substantial effect on language-model performance in downstream tasks. [^1] Code available here.
Parity-Aware Byte-Pair Encoding:
Improving Cross-lingual Fairness in Tokenization
Negar Foroutan1thanks: Equal contribution, †Equal supervision.   Clara Meister211footnotemark: 1   Debjit Paul1 Joel Niklaus3   Sina Ahmadi4   Antoine Bosselut1   Rico Sennrich4 1EPFL  2ETH Zurich  3Niklaus.ai  4University of Zurich
1 Introduction
At a time of rapid innovation and constant change in natural language processing (NLP), tokenization continues to be a foundational and comparatively stable component of NLP pipelines. Tokenization is the transformation of raw sequences of bytes [^2] Early work considered characters the “base unit” of strings, but bytes have become popular because their fixed 256-symbol vocabulary can encode any character from any encoding, eliminating out-of-vocabulary issues. into sequences of byte-spans, i.e., subwords; it enables computational efficiency and provides essential inductive biases by defining meaningful textual units. This design choice can have a major impact on various aspects of model performance (Bostrom and Durrett, 2020; Ali et al., 2024; Goldman et al., 2024).
The predominant tokenization algorithms—Byte Pair Encoding (BPE; Sennrich et al., 2016) and UnigramLM (Kudo, 2018), for example—select the vocabulary by maximizing frequency-based objectives computed over an entire training corpus. In multilingual corpora, this global criterion inevitably favors the languages with the greatest representation. Under vocabulary size constraints, subwords that primarily benefit high-resource languages are preferentially included, often at the expense of those needed for lower-resource languages. This bias has both qualitative and economic consequences. NLP models trained on fragmented or semantically incoherent tokenizations lose valuable inductive biases and tend to perform worse. At the same time, texts in lower-resource languages—often tokenized into more tokens—incur higher computational costs from language-model-based services charging based on token count, which disproportionately burdens users of underrepresented languages and exacerbates existing inequalities.
In the effort to mitigate these inequities, we introduce Parity-aware BPE. The classic version of BPE learns its vocabulary by repeatedly selecting the subword pair with the highest corpus-level co-occurrence count; it adds the concatenation of this pair to the vocabulary and replaces all pair co-occurrences with the new symbol. [^3] This iterative merging process can be viewed as a form of data compression, where frequent subword sequences are replaced with shorter representations. Parity-aware BPE is a simple variant of this algorithm, retaining the iterative framework but redefining the merge selection rule: at each step, it computes co-occurrence statistics separately for each language and then uses statistics from the language with the current worst compression rate for selecting the next merge. In other words, instead of greedily maximizing a global objective, Parity-aware BPE performs a “fair-max” update that targets the worst-off language. It thus progressively equalizes string compression rates across languages.
Empirically, we find that Parity-aware BPE leads to better token-count parity across languages compared to Classical BPE while maintaining comparable global compression rates. Evaluations on 13 multilingual benchmarks show that models trained with a Parity-aware tokenizer match or exceed downstream performance compared to those trained with a Classical BPE tokenizer. Fairness metrics improve significantly, indicating a more even distribution of token costs without sacrificing efficiency or downstream performance. In short, Parity-aware BPE narrows existing tokenizer-induced disparities between languages, ensuring more equitable resource allocation and balanced performance across diverse linguistic use cases.
2 Text Tokenization
Text can be decomposed at many granularities: graphemes, Unicode code points, or multi-character tokens; but at the most fundamental digital layer, every string is represented as a sequence of bytes, the foundation on which all other units are constructed. Let $b\in\mathcal{B}=\{0,\dots,255\}$ denote an individual byte. A finite byte-string is written $\mathbf{b}\in\mathcal{B}^{*}$, where $\mathbf{b}=b_{1}b_{2}\cdots b_{|\mathbf{b}|}$. Throughout, bytes are treated as atomic symbols. Note that all subsequent definitions apply if one substitutes bytes with another finite alphabet of atomic units.
2.1 Byte-level Tokenizers
[…]
  • Vocabulary $\mathcal{V}\subset\mathcal{B}^{+}$: a finite set of non-empty byte-spans, often called subwords.
  • Tokenization function $\tau:\mathcal{B}^{*}\!\to\!\mathcal{V}^{*}$: a mapping from byte-strings $\mathbf{b}$ to sequences of tokens $\mathbf{v}=v_{1},v_{2},\ldots$.
[…]
Pre-tokenization and Normalization.
Many tokenization algorithms include a pre-tokenization (and often a normalization) step that segments or rewrites raw byte strings according to deterministic criteria. Pre-tokenization can encompass several operations, including Unicode normalization or splitting on whitespace. Notably, pre-tokenization determines subword boundaries, and thus also determines the set of possible candidates for the vocabulary as well as the attainable compression rate. For example, if whitespace is used as a subword boundary, languages without explicit whitespace (e.g., Chinese, Japanese) or with rich morphology may have the potential for higher compression. For simplicity, we assume this step is baked into $\tau$.
2.2 Text Compression
Why does mapping a raw byte-string to a sequence of larger subword tokens help a language model (LM)? The precise inductive biases this process imbues remain an open research question (Zouhar et al., 2023a; Schmidt et al., 2024), but one plausible explanation is the compression it provides: a good tokenizer tends to map each input $\mathbf{b}$ to a shorter sequence of tokens, reducing the length of the model’s effective input and, potentially making learning easier. At the very least, it can significantly reduce model-side computations.
[…]
$\mathrm{CR}(\mathbf{b};\tau)\;\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\;\frac{\lvert\mathbf{b}\rvert_{u}}{\lvert\tau(\mathbf{b})\rvert}$ (2)
where $|\mathbf{b}|_{u}$ denotes the length of $\mathbf{b}$ in terms of a given normalization unit $u$ (e.g., characters, words, lines, or simply bytes). In words, $\mathrm{CR}(\mathbf{b};\tau)$ measures the multiplicative factor by which our original sequence length is reduced after tokenization. A higher $\mathrm{CR}$ indicates stronger compression.
We are generally interested in a tokenizer’s average compression, which can be estimated from a corpus $\mathcal{D}$:
$\mathrm{CR}(\mathcal{D};\tau)\;\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\;\frac{1}{\lvert\mathcal{D}\rvert}\sum_{\mathbf{b}\in\mathcal{D}}\frac{\lvert\mathbf{b}\rvert_{u}}{\lvert\tau(\mathbf{b})\rvert}$ (3)
This quantity provides an estimate of the average number of tokens that an autoregressive LM must process per unit $u$ of raw text. Tokenizers differ in how small they can make $\mathrm{CR}(\mathcal{D};\tau)$ while remaining lossless. Further, this rate can vary across strings from different languages, which motivates our last definition: language-specific compression rate.
Let $\mathcal{L}=\{\ell^{(r)}\}_{r=1}^{R}$ be our set of languages and $\mathcal{M}=\{(\mathbf{b}^{(s)},\ell^{(s)})\}_{s=1}^{S}$ be a labeled multilingual corpus, i.e., a corpus where each byte-string $\mathbf{b}^{(s)}$ is labeled with its respective language. For fixed $T$, we define a language’s compression rate as
$\mathrm{CR}(\ell;\tau)\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\mathrm{CR}(\mathcal{D}_{\ell};\tau)$ (4)
where $\mathcal{D}_{\ell}=\{\mathbf{b}^{(s)}:(\mathbf{b}^{(s)},\ell^{(s)})\in\mathcal{M},\,\ell^{(s)}=\ell\}$.
Compression Rate as a Notion of Fairness.
Many commercial technology services offer APIs that bill per token; a service’s processing speed also scale with the number of tokens in the input and output. Token counts thus dictate both the economics and the latency of these services. For a given byte sequence $\mathbf{b}$, the number of tokens produced by tokenization function $\tau$ is determined by the tokenizer’s compression rate $\mathrm{CR}(\mathbf{b};\tau)$. $\mathrm{CR}(\ell;\tau)$—the expected compression rate over sequences from language $\ell$—is thus a direct proxy for the average per-language cost (and expected latency) of using one of these services. Variance in $\mathrm{CR}(\ell;\tau)$ across languages thus implies different user costs purely as a function of language choice (Petrov et al., 2023), and whether $\mathrm{CR}(\ell;\tau)$ is comparable across languages is therefore one way of measuring tokenizer fairness. Byte Pair Encoding, discussed next, optimizes for compression rate across a corpus without regard for language. Our subsequent adjustment to Classical BPE adds an auxiliary objective of equalizing $\mathrm{CR}(\ell;\tau)$ across languages.
2.3 Byte Pair Encoding
Byte Pair Encoding (BPE; Sennrich et al., 2016) is one popular algorithm for creating a tokenizer adapted from the byte-pair compression scheme of Gage (1994). In short, BPE tokenizes text by iteratively merging adjacent tokens whose token-types (i.e., subwords) were observed to co-occur frequently in the training data.
The notion of a merge lets us formalize this procedure. A merge is defined as an ordered pair $m=(v,v^{\prime})$ with $v,v^{\prime}\in\mathcal{V}$. The application of a merge to a token sequence replaces every bigram token $v,v^{\prime}$ by a single token $v\circ v^{\prime}$. Each bigram token replacement shortens the token sequence by exactly one token, thereby compressing the sequence. To tokenize a piece of text with a BPE tokenizer, we start from its representation as a byte-string, i.e., a sequence of base bytes—all of which necessarily appear in our tokenizer’s vocabulary. We then iteratively apply a given list of merges to that sequence. Note that because the merge list is fixed in advance, the encoding is deterministic. Intuition for the merge procedure is perhaps best acquired by a small example:
Example 2.1 (Example of the iterative application of merge sequence $\mathbf{m}$ to byte sequence $\mathbf{b}$ ).
[…]
$\text{Step 2}:\texttt{ba b}\!\to\!\texttt{bab}\,\,\Longrightarrow\,\,\texttt{ba bab}$
[…]
  • $\mathcal{V}=\mathcal{B}\cup\{v\circ v^{\prime}:(v,v^{\prime})\in\mathbf{m}\}$
  • $\tau_{\mathbf{m}}$ carries out the procedure described above, i.e., it applies each $m_{k}$ to an input byte-string $\mathbf{b}$ in the prescribed order. In example˜2.1, $\tau_{\mathbf{m}}(\mathbf{b})={\texttt{ba},\texttt{bab}}$.
[…]
We use the $\mathbf{m}$ subscript here to make explicit the tokenization function’s dependence on $\mathbf{m}$.
Learning $\mathbf{m}$.
The BPE algorithm seeks the merge list $\mathbf{m}^{*}$ (subject to a size constraint $K$) that maximizes the compression rate of the given corpus:
$\mathbf{m}^{*}\;=\;\max_{\mathbf{m}:|\mathbf{m}|=K}\,\mathrm{CR}(\mathcal{D};\tau_{\mathbf{m}})$ (5)
BPE takes a greedy approach to choosing $\mathbf{m}$, finding an approximate solution to eq.˜5 (Zouhar et al., 2023b). It starts with the singleton-byte vocabulary $\mathcal{V}_{0}=\mathcal{B}$ and repeatedly greedily enlarges the vocabulary. At each of $K$ steps, the current tokenizer $\tau_{\mathbf{m}_{<k}}$ is applied to the entire training corpus, and the algorithm counts how often every adjacent pair of tokens occurs. The subword-type pair with the highest count, which we denote as $(v^{\star},v^{\prime\star})$, is deemed the most “compressive.” Its concatenation $v^{\star}\circ v^{\prime\star}$ is added to the vocabulary, the merge $m_{k}=(v^{\star},v^{\prime\star})$ is recorded, and every occurrence of the bigram $(v^{\star},v^{\prime\star})$ in the corpus is replaced by the new token so the next iteration works with updated token sequences. Repeating this process $K$ times yields the ordered list $\mathbf{m}=[m_{1},\ldots,m_{K}]$ and the final vocabulary $\mathcal{V}_{K}$. When encoding a new text, $\tau_{\mathbf{m}}$ simply applies these merges in the same order. The pseudocode for the algorithm is provided in Alg.˜1 in App.˜A.
3 Parity-aware Byte Pair Encoding
Classical BPE chooses merges that maximize a global frequency objective, implicitly favoring the compression of languages with a larger presence in the training corpus. Here we introduce Parity-aware BPE, which replaces this global objective with a max–min criterion: at every step, it selects the merge that most improves the language currently suffering the poorest compression rate. In this section, we formalize the objective and describe the resulting algorithm.
3.1 Greedy min-max objective
Our adjustment to the Classical BPE objective (eq.˜5) explicitly encodes our earlier notion of tokenizer fairness: equality across per-language compression rates. Formally, parity-aware BPE seeks a merge list $\mathbf{m}=[m_{1},\dots,m_{K}]$ that maximizes the minimum compression rate across languages:
$\mathbf{m}^{\star}=\max_{\mathbf{m}:|\mathbf{m}|=K}\,\min_{\ell}\,\mathrm{CR}(\ell;\tau_{\mathbf{m}})$ (6)
This min-max objective trades a small amount of global compression for fairness across languages.
3.2 Algorithm
Parity-aware BPE retains the greedy iterative framework of Classical BPE but changes which statistics are inspected each time a merge is added. At merge step $k$, it identifies the language with the worst compression under the tokenizer defined by the merge list thus far ($\mathbf{m}_{<k})$
$\ell^{\star}\;=\;\arg\min_{\ell\in\mathcal{L}}\mathrm{CR}(\ell;\tau_{\mathbf{m}_{<k}})$ (7)
To choose the next merge, it uses the same maximum pair count criterion as Classical BPE, albeit with pair counts computed over only $\mathcal{D}_{\ell^{\star}}$—the portion of the corpus corresponding to $\ell^{\star}$. The rest of the algorithm follows the Classical BPE procedure: the chosen merge is applied to all texts (i.e., across $\mathcal{D}_{\ell}$ $\forall\ell$) [^4] Crucially, this is what distinguishes our algorithm from a combination of monolingual merge lists (Petrov et al., 2023), allowing us to find more “compressive” merges. and the procedure is repeated for $k=1,\dots,K$, yielding the final merge list $\mathbf{m}$. We provide pseudocode in Alg.˜2.
Cross-lingual Compression Rate Comparison.
Parity-aware BPE relies on the comparison of $\mathrm{CR}(\ell;\tau_{\mathbf{m}})$ across different $\ell$. The choice of normalization unit $u$ has a large impact on the measured $\mathrm{CR}(\ell;\tau_{\mathbf{m}})$ and even when $u$ is held constant across measurements for different languages, if not considered carefully, the choice can introduce bias into the comparison. As concrete examples, certain normalization units are more appropriate in some languages than in others, e.g., whitespace-delimited “words” are ill-defined in many languages; although principled and universal, even normalizing by byte can skew perceived compression because scripts differ greatly in average bytes per character (e.g., ASCII vs. UTF-8 CJK). Parallel corpora provide a principled solution: computing compression rates over aligned segments (sentences, lines, or documents) normalizes by content rather than script, making cross-language comparisons more meaningful. We therefore recommend the use of a parallel corpus for computing eq.˜7. Notably, this evaluation corpus need not be the same one used for computing subword pair frequency statistics, for which a larger corpus with only language annotation is necessary. For generality, we thus differentiate between the corpora used to compute frequency statistics and in computing eq.˜7, referring to them as our training and development datasets, respectively. Alg.˜2 makes this difference explicit. We present experimental results both with a separate, parallel development set and using a single (not parallel) multilingual dataset for all computations.
Complexity and Data Requirements.
Relative to Classical BPE, Parity-aware BPE incurs only a $O(|\mathcal{L}|)$ overhead per-merge from recomputing the language-specific compression rates on the dev set. Parity-aware BPE retains the same asymptotic complexity as Classical BPE, requiring only some modest additional bookkeeping. The need for a parallel multilingual corpus can at first seem prohibitive, but several pragmatic design choices can reduce the burden of this requirement. A small, aligned dataset suffices to drive the max-min decision in eq.˜7, and the training dataset need not have this level of annotation. In addition, automatic language ID tools or script heuristics can help provide language labels when none are readily available.
Also note that only the BPE learning phase differs; there is no algorithmic change to the tokenization function itself.
3.3 Algorithmic Variants
Preliminary experiments have shown several challenges with parity-aware BPE, which we address by introducing two variants.
Hybrid parity-aware BPE.
Model developers may want to include and tokenize data for which parallel data is not available or where this concept does not even apply, such as programming code. Also, they may not want to guarantee full parity, but still give a high weight to global compression. We support these goals with a hybrid learning algorithm that uses the global objective of vanilla BPE (eq.˜5) for the first $K$ merges, then switches to the parity-aware objective (eq.˜6) for another $J$ merges. $K$ and $J$ can be chosen by model developers to assign budgets to global compression and fairness.
Moving-window balancing.
There may be a point where the compression in a language no longer or barely improves, even if it is repeatedly chosen for the next merge. This could happen if the development dataset (the dataset used to choose the language) is too small or does not match the domain or language variant of the training dataset, [^5] Kreutzer et al. (2022) discuss possible quality issues such as wrong or ambiguous language codes. or if $\lvert\tau(\mathbf{b})\rvert$ approaches the length of the pre-tokenized sequence. To prevent our algorithm from being “stuck” selecting the same language exclusively, we track the $W$ most recent languages selected in eq.˜7, and do not select a language if it occurs more than $\alpha\frac{W}{|\mathcal{L}|}$ times in this moving window.
4 Experimental Setup
We conduct experiments to evaluate the effectiveness of Parity-aware BPE, comparing it against baseline tokenization methods: Classical BPE and UnigramLM. All tokenizers are byte-level.
4.1 Tokenizer Training
Training Data.
We train tokenizers using the multilingual C4 (mC4) corpus (Xue et al., 2021; Raffel et al., 2020). [^6] https://huggingface.co/datasets/allenai/c4 For choosing the focus language at each merge step when training Parity-aware BPE tokenizers (i.e., computing eq.˜7), we use the dev portion of FLORES+ (NLLB Team et al., 2024) as our multilingual development corpus—except for the no-dev systems, for which the training corpus is used to measure compression rate with bytes as normalization unit. To investigate how the number of languages, their linguistic diversity, and the variety of writing systems influence tokenizers, we consider two language sets: one with 30 languages (30-lang) and another with 60 languages (60-lang). For each of these language sets, we create two dataset versions: one with uniform quantities of data per language (balanced) and one with per-language quantities proportional to amounts in the mC4 dataset (unbalanced). We present results for the unbalanced datasets here, as this is arguably the more realistic setting, with results for the balanced setting shown in Fig.˜3. To enable tokenizer analyses as a function of different dataset qualities, we categorize the languages in each set based on the amount of training data available and the script family, performing some of our analyses by these categories. Languages with $1\text{M}$ examples are considered high-resource; those with $500\text{k}-1\text{M}$ examples are medium-resource; and those with fewer than $500\text{k}$ examples are classified as low-resource. The full list of languages included in each set and the script family groupings are presented in Table˜6.
Hyperparameter Settings.
We look at vocabulary sizes $128k$ and $256k$. For hybrid systems, we learn half of the merges using the global strategy, and the second half using the parity-aware strategy. For systems with moving-window balancing (window), we use a window size of 100, and $\alpha=2$.
4.2 Evaluation
Our evaluations consist of task-independent tokenizer properties (intrinsic metrics) and downstream model performance (extrinsic metrics).
4.2.1 Intrinsic Metrics
We measure a variety of intrinsic tokenizer metrics on the devtest portion of FLORES+. These metrics encompass basic tokenization properties, information-theoretic measures, cross-linguistic fairness, and morphological alignment. All metrics are computed both globally and per-language to capture language-specific tokenization behavior. Normalization units differ across metrics, both in the effort to control for confounding factors and to tailor the metric to the tokenizer quality it is trying to measure. For example, morphologically motivated units can better reflect linguistic structure, and using character or document–level units can partially normalize the large differences in average bytes-per-character observed across writing systems (e.g., Latin vs. UTF-8–encoded CJK [^7] Chinese, Japanese and Korean scripts. scripts). For the sake of space, we provide brief metric descriptions here; more detailed definitions for all metrics, including formulae and implementation details, can be found in App.˜B.
  • Fertility measures the average number of tokens produced per normalization unit by a tokenizer; whitespace-delimited words are often the unit of interest (and are the units used in our computations). In this case, fertility quantifies how many tokens (on average) a word is broken up into.
  • Compression Rate (CR) (as defined in §˜2) is a measure of the degree to which a unit of text has been shrunk after applying the given tokenizer (higher is better). Because we evaluate on parallel corpora, we can use documents as the normalization unit to control for differences in scripts’ average bytes-per-character.
  • Vocabulary Utilization is the fraction of the tokenizer’s vocabulary that actually appears in the evaluation corpus. Low utilization for a language signals wasted capacity or—when there are large differences across languages—biased vocabulary allocation.
  • Tokenizer Fairness Gini adapts the Gini coefficient to the per-language tokenization cost distribution (e.g., tokens per line (document) in a parallel corpus). Values near 0 mean equal cost across languages; values closer to 1 indicate inequality.
  • MorphScore (Arnett et al., 2025) measures how well token boundaries align with true morpheme boundaries, computed as morpheme-level precision/recall (and F1). High scores mean tokens respect morphological structure; low precision implies over-segmentation, while low recall may suggest under-segmentation.
For completeness, we also track Type–Token Ratio and Average Token Rank (vocabulary diversity; Limisiewicz et al., 2023) as well as Rényi entropies (distributional concentration; Zouhar et al., 2023a).
4.2.2 Extrinsic Metrics.
For extrinsic evaluation, we train models using different tokenizers and assess their performance across a range of downstream tasks.
Model Architecture and Pretraining Data.
We train decoder-only Transformer models Vaswani et al. (2017) following the LLaMA architecture Touvron et al. (2023) with 3 billion (33B) parameters. Full details on model configurations and training parameters are provided in App.˜D. Models are trained on the FineWeb2 corpus (Penedo et al., 2025). We adopt temperature sampling with $\tau=3.3$, following recommendations from prior work (Raffel et al., 2020; Conneau et al., 2020). We use the total 100B tokens to train each model.
Benchmarks.
We evaluate the models using perplexity on a held-out validation set from the respective pretraining datasets. In addition, we assess downstream performance on a suite of multilingual benchmarks. Results are aggregated per language to produce a score for each model-language pair. A full list of benchmarks and aggregation procedures is provided in the App.˜E.
5 Results and Analysis
We present and analyze results on both intrinsic and extrinsic metrics. Within a set of comparisons, we fix the data distribution (balanced or unbalanced) and vocabulary size ($128k$ or $256k$).
5.1 Intrinsic Evaluation
Figure 1: Vocabulary utilization for $128k$ tokenizers on the unbalanced 30-lang grouped by dataset resource levels. -- lines indicate macro-averages across groups.
Tokenizer Type-Token Ratio Fertility Compression Rate Rényi Entropy ($\alpha$=2.5) Gini Coefficient MorphScore Precision MorphScore Recall Classical 0.0743 4.260 $\pm$ 0.049 0.0303 $\pm$ 0.0001 8.13 0.064 0.412 $\pm$ 0.051 0.456 $\pm$ 0.049 UnigramLM 0.0475 4.612 $\pm$ 0.042 0.0228 $\pm$ 0.0001 4.68 0.094 0.153 $\pm$ 0.037 0.268 $\pm$ 0.053 Parity-aware 0.0765 4.204 $\pm$ 0.049 0.0300 $\pm$ 0.0001 8.12 0.011 0.407 $\pm$ 0.051 0.457 $\pm$ 0.049 Parity-aware (hybrid) 0.0770 4.191 $\pm$ 0.049 0.0303 $\pm$ 0.0001 8.10 0.018 0.412 $\pm$ 0.051 0.457 $\pm$ 0.049 Parity-aware (window) 0.0788 4.219 $\pm$ 0.050 0.0302 $\pm$ 0.0001 8.11 0.013 0.405 $\pm$ 0.049 0.453 $\pm$ 0.047 Parity-aware (window+hybrid) 0.0794 4.203 $\pm$ 0.050 0.0305 $\pm$ 0.0001 8.09 0.022 0.416 $\pm$ 0.049 0.460 $\pm$ 0.047 Parity-aware (no-dev) 0.0772 4.310 $\pm$ 0.050 0.0303 $\pm$ 0.0001 8.12 0.059 0.423 $\pm$ 0.051 0.466 $\pm$ 0.049
Table 1: Intrinsic evaluation of $128k$ tokenizers on the (unbalanced) 30-lang dataset. Values are global statistics across the parallel corpus, except for MorphScore, which is macro-averaged across available languages.
Results in Table˜1 show that all variants of Parity-aware BPE outperform Classical BPE in terms of the Gini coefficient, indicating more equitable token costs per document across languages. Among these variants, the base Parity-aware BPE emerges as the “fairest” tokenizer. Notably, Classical BPE and the Parity-aware BPE variants attain almost identical compression and Rényi entropies; we take this as evidence that the parity-aware variants match global efficiency while redistributing it more evenly. In addition, Parity-aware variants reduce fertility and increase MorphScore and vocabulary utilization, signaling better alignment with morphological boundaries and fairer vocabulary allocation. As could perhaps be expected, the no-dev variant of Parity-aware BPE performs most similarly to Classical BPE, closing only part of the Gini gap but matching it almost exactly on every other metric. This observation demonstrates the importance of thoughtfully consideration of normalization units for the effectiveness of the algorithm; future work could address this by explicitly compensating for differences in cross-language length statistic, e.g., by introducing a per-language multiplicative factor. [^8] For example, in the absence of a multi-parallel corpus, developers could estimate the desired compression rate from a number of parallel corpora. By contrast, the hybrid and window variants land between Parity-aware and Classical BPE on many metrics: they recover a small slice of global compression while reducing the Gini inequity coefficient of Classical BPE by roughly three-quarters and achieving the lowest fertility of all runs. Taken together, the variants outline a smooth fairness–efficiency frontier, allowing practitioners to select the point that best suits their resource constraints and fairness targets.
Language Classical BPE
Parity-aware
Parity-aware
(window+hybrid)
Random
Arabic 38.19 $\pm$ 2.90 39.04 $\pm$ 2.89 38.84 $\pm$ 2.90 32.00
Bengali 24.95 $\pm$ 3.09 23.54 $\pm$ 2.98 23.91 $\pm$ 3.01 25.00
German 32.92 $\pm$ 3.14 34.78 $\pm$ 3.66 36.82 $\pm$ 4.04 30.62
Greek 41.95 $\pm$ 3.18 42.55 $\pm$ 3.22 43.16 $\pm$ 3.25 37.50
Spanish 37.53 $\pm$ 2.66 38.83 $\pm$ 2.71 39.30 $\pm$ 2.75 32.77
Persian 42.80 $\pm$ 5.39 39.15 $\pm$ 5.27 39.15 $\pm$ 5.27 25.00
French 38.67 $\pm$ 3.90 36.59 $\pm$ 2.84 37.10 $\pm$ 2.82 32.00
Hindi 33.92 $\pm$ 2.25 33.92 $\pm$ 2.24 33.86 $\pm$ 2.24 30.62
Indonesian 38.95 $\pm$ 2.62 40.55 $\pm$ 2.66 40.46 $\pm$ 2.66 35.00
Italian 32.82 $\pm$ 2.86 35.01 $\pm$ 3.00 34.62 $\pm$ 2.98 27.22
Japanese 37.43 $\pm$ 2.39 37.45 $\pm$ 2.39 37.43 $\pm$ 2.39 34.00
Korean 33.00 $\pm$ 5.22 33.00 $\pm$ 5.22 34.33 $\pm$ 5.29 25.00
Polish 29.75 $\pm$ 2.50 31.14 $\pm$ 2.60 28.97 $\pm$ 2.49 23.75
Portuguese 33.63 $\pm$ 2.81 33.15 $\pm$ 2.77 33.06 $\pm$ 2.77 27.50
Russian 36.36 $\pm$ 2.27 36.21 $\pm$ 2.26 36.57 $\pm$ 2.28 32.77
Tamil 31.32 $\pm$ 2.81 32.25 $\pm$ 2.90 32.19 $\pm$ 2.90 31.25
Telugu 32.73 $\pm$ 2.61 33.52 $\pm$ 2.61 33.26 $\pm$ 2.61 30.00
Turkish 39.04 $\pm$ 2.89 38.46 $\pm$ 2.83 37.89 $\pm$ 2.75 35.00
Vietnamese 33.69 $\pm$ 2.31 33.87 $\pm$ 2.27 33.80 $\pm$ 2.30 29.50
Chinese 38.43 $\pm$ 2.11 38.58 $\pm$ 2.11 38.32 $\pm$ 2.10 35.00
English 43.04 $\pm$ 1.84 44.15 $\pm$ 1.85 43.74 $\pm$ 1.85 35.50
Thai 40.76 $\pm$ 1.62 40.96 $\pm$ 1.63 41.06 $\pm$ 1.63 37.50
Table 2: Average LM performance (accuracy %) across 13 multilingual benchmarks (tokenizers trained on the unbalanced 30-lang dataset with $128k$ vocab size). The Random column shows the expected accuracy of a random classifier. Best performance per language is bolded. Benchmark descriptions and details for each language are provided in ˜E and 7, respectively.
Fig.˜1 presents vocabulary utilization grouped by resource tier, as defined in §˜4.1. For high-resource languages, the results indicate that parity-aware BPE (no-dev) performs comparably to Classical BPE in terms of vocabulary utilization, while the other variants provide worse vocabulary utilization. In contrast, for low- and medium-resource languages, the hybrid and window variants achieve higher vocabulary utilization, highlighting their effectiveness at providing fairer vocabulary allocation across languages. We show per-language results for compression rate and vocabulary utilization in Fig.˜2 and 3. We observe that Parity-aware tokenizers attain substantially more uniform compression across languages. While they also generally achieve higher vocabulary utilization rates across most languages, the level of benefit varies by language.
Vocabulary Size.
Repeating the experiment with a $256k$ vocabulary (Table˜5 and Fig.˜4) yields the same conclusion: parity-aware BPE tokenizers consistently outperform Classical BPE in terms of metrics indicative of cross-lingual fairness with minimal changes to more performance-oriented measures of tokenizer quality.
5.2 Extrinsic Evaluation
Table˜2 presents the performance of LMs trained with three different $128k$ tokenizers: Classical BPE, Hybrid Parity-aware BPE, and Hybrid Parity-aware BPE (moving-window), evaluated on the 30-lang set. For each language, we report mean performance, standard errors, and a random baseline to account for varying benchmark counts (Table˜7). The evaluation assesses Parity-aware BPE’s impact on downstream performance, particularly in languages where Classical BPE is efficient. Results indicate that Parity-aware BPE maintains performance across languages: accuracy changes relative to Classical BPE are small. Models trained with the hybrid variant show a median per-language change in accuracy of +0.19+0.19 percentage points, with 14 languages improving and 6 declining; the window+hybrid leads to very similar changes in accuracy. These results confirm that parity-aware tokenizers can handle diverse languages without compromising LM performance. We show per-language perplexity results in Fig.˜6 in Fig.˜3. Here, we see that models trained with parity-aware tokenizers show much more uniform perplexity across languages, whereas Classical BPE yields a handful of languages with markedly higher perplexity.
6 Related Work
Multilingual Tokenization.
Despite their popularity, BPE and similar subword tokenization algorithms often underperform in multilingual settings due to limited handling of spelling variation and morphological complexity (Bostrom and Durrett, 2020). Key metrics like tokenization parity and fertility directly impact computational costs and model performance. Previous work has examined vocabulary allocation strategies: Zhang et al. (2022) find that increasing vocabulary size enhances NMT robustness across different scripts, while Gowda and May (2020) show that BPE merges can be tuned to address sequence length issues. Rust et al. (2021) find that specialized monolingual tokenizers integrated into multilingual systems can improve performance; however, recent evidence suggests that the optimal vocabulary size varies with the task and model (Dagan et al., 2024). In terms of multilingual vocabulary construction, Chung et al. (2020) explore clustering-based sharing of subword units across languages, and Limisiewicz et al. (2023) propose an explicit tokenizer-merging algorithm to combine vocabularies of separate per-language tokenizers. Tokenization-free models like canine (Clark et al., 2022) and ByT5 (Xue et al., 2022) also offer a potential route forward for better handling of multilingual data.
Tokenization Bias and Recent Advances.
Recent research highlights biases from tokenization in LLMs. While Wan (2022) argues that character- and byte-level representations are intrinsically fair, [^9] They report more random performance with these representations, which we do not view as fairness. other studies (Petrov et al., 2023; Ahia et al., 2023) show that tokenization differences across languages; even at character and byte levels; affect costs, latency, and contextual understanding. This has spurred efforts like Aya Aryabumi et al. (2024) and methods to mitigate tokenization unfairness (Fujii et al., 2024; Abboud and Oz, 2024; Limisiewicz et al., 2024). Although newer character- and byte-level models use compression techniques such as entropy-based patching Pagnoni et al. (2025), cross-lingual parity of these representations remains unstudied. Finally, Ali et al. (2024) found parity metrics weakly predict LLM performance, but their results were confounded by differing tokenizers and vocabularies, unlike our algorithms, which improve parity under a fixed vocabulary.
7 Discussion and Conclusion
Tokenizers optimized using standard algorithms and data can lead to disparities in users’ costs and experiences as a result of their choice of language. Parity-optimized tokenization can remedy this by explicitly balancing compression across languages, enabling fairer treatment of users of low-resource languages. Parity-aware BPE implements this idea: it was designed to improve cross-lingual tokenization parity, and our experiments confirm that it does. On the unbalanced 30-language set, the Gini coefficient of per-line token costs falls from $0.064$ with Classical BPE to $0.011$ with our parity-aware variant while compression ratios for most variants stay competitive with Classical BPE, often improving when looking at averages across languages on the whole. Crucially, this fairness gain does not come at the expense of downstream quality: across 13 multilingual benchmarks, models trained with parity-aware tokenizers either outperform or stay within a single standard error of the Classical BPE baseline for every tested language (Table˜2).
Overall, the trade-offs required for using parity-aware BPE are minimal. During the learning stage, the algorithm adds only an $\mathcal{O}(\mathcal{L})$ pass per merge for recomputing language-level compression rates on a dev corpus, leaving the asymptotic complexity identical to Classical BPE. Further, we observe empirically that a small, sentence-aligned development set is sufficient to drive the “fair-max” decision. When resource or domain mismatches make full equality undesirable, hybrid and moving-window variants further let practitioners trade off global compression versus strict parity; our empirical results validate that these variants perform well in practice. From the model-developer’s perspective, parity-aware BPE is a drop-in replacement: it requires no architectural changes and minimal changes to the tokenizer pipeline.
Making the tokenization step of the NLP pipeline more equitable is therefore not just desirable but feasible. Parity-aware BPE offers a clear avenue towards this goal by building fairness into the tokenizer itself. It achieves this via only a simple modification to Classical BPE—choosing the merge that most benefits the currently worst-compressed language. Yet with this modification, parity-aware BPE dramatically narrows token-count disparities, mitigating the hidden “token tax” imposed on speakers of low-resource languages. It does so without compromising overall compression or downstream task accuracy. Future work can push this agenda further by extending parity objectives to alternative tokenization schemes (e.g., UnigramLM, WordPiece) and other modalities such as speech and vision, as well as by developing benchmarks and metrics for fairness assessments in tokenization beyond compression parity.
Limitations
Our study uses parallel corpora to estimate per-language costs; in domains where aligned documents are unavailable or difficult to obtain, using unaligned corpora and alternative normalization units for making the language choice may introduce bias. While we consider 60 languages and two vocabulary sizes, the interplay between tokenization parity and model scaling still needs to be explored for much larger models and for code or multimodal inputs. Finally, fairness here is defined purely in terms of token counts. While we measure other potential quantification of fairness (e.g., morphological alignment), there are still other notions that are unaccounted for. We leave optimization for these metrics during tokenizer learning to future work.
Acknowledgements
RS acknowledges support by the Swiss National Science Foundation through the MUTAMUR project (no. 213976). We also gratefully acknowledge the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, a Meta LLM Evaluation Research Grant, and the Swiss AI Initiative through a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a06 on Alps.
Appendix A Pseudocode
Input: Corpus $\mathcal{D}$; number of merges $K$
Output: Vocabulary $\mathcal{V}_{K}$; merge sequence $\mathbf{m}_{K}$
1
[…]
3 $\mathbf{m}_{0}\leftarrow\langle\,\rangle$
4 for $k\leftarrow 1$ to $K$ do
 // Count all adjacent token pairs
5 $\textit{Pairs}\leftarrow\{\}$
6 foreach occurrence of consecutive tokens $v\,v^{\prime}$ in $\mathcal{D}$ where $v,v^{\prime}\in\mathcal{V}_{k-1}$ do
7    $\textit{Pairs}[(v,v^{\prime})]\leftarrow\textit{Pairs}[(v,v^{\prime})]+1$
8
9
10 $(v^{\star},v^{\prime\star})\leftarrow\arg\max_{(v,v^{\prime})}\textit{Pairs}[(v,v^{\prime})]$
11
12 $w^{\star}\leftarrow v^{\star}\!\circ\!v^{\prime\star}$
13
 // Update vocabulary and merge sequence
14 $\mathcal{V}_{k}\leftarrow\mathcal{V}_{k-1}\cup\{w^{\star}\}$
15 $\mathbf{m}_{k}\leftarrow\mathbf{m}_{k-1}+\langle(v^{\star},v^{\prime\star})\rangle$
16
 // Replace all occurrences in corpus
17 foreach occurrence of $v^{\star}\,v^{\prime\star}$ in $\mathcal{D}$ do
18      Replace $v^{\star}\,v^{\prime\star}$ with $w^{\star}$
19
20
21
return $\mathcal{V}_{K}$, $\mathbf{m}_{K}$
Algorithm 1 Algorithm for learning $\mathbf{m}$ using Classical BPE.
Input: $\bigl{\{}\mathcal{D}_{\ell}\bigr{\}}_{\ell\in\mathcal{L}}$ (multilingual training corpus);
[…]
    $K$ (number of merges)
Output: $\mathcal{V}_{K}$ (vocabulary); $\mathbf{m}_{K}$ (merge list)
1 $\mathcal{V}_{0}\leftarrow\mathcal{B}$;    $\mathbf{m}_{0}\leftarrow\langle\,\rangle$
2 for $k\leftarrow 1$ to $K$ do
 // Calculate compression rate for each language
3 foreach language $\ell\in\mathcal{L}$ do
[…]
5
[…]
7
 // Consider token pairs only in $\mathcal{D}_{\ell^{\star}}$
8 $\textit{Pairs}\leftarrow\{\}$
9 foreach occurrence of consecutive tokens $v\,v^{\prime}$ in $\mathcal{D}_{\ell^{\star}}$ where $v,v^{\prime}\in\mathcal{V}_{k-1}$ do
10    $\textit{Pairs}[(v,v^{\prime})]\leftarrow\textit{Pairs}[(v,v^{\prime})]+1$
11
12 $(v^{\star},v^{\prime\star})\leftarrow\arg\max_{(v,v^{\prime})}\textit{Pairs}[(v,v^{\prime})]$
13 $w^{\star}\leftarrow v^{\star}\!\circ\!v^{\prime\star}$
14
 // Update vocabulary and merge list
15 $\mathcal{V}_{k}\leftarrow\mathcal{V}_{k-1}\cup\{w^{\star}\}$
16 $\mathbf{m}_{k}\leftarrow\mathbf{m}_{<k}+\langle(v^{\star},v^{\prime\star})\rangle$
17
 // Apply merge across all languages
18
19 foreach language $\ell\in\mathcal{L}$ do
[…]
21         Replace $v^{\star}\,v^{\prime\star}$ with $w^{\star}$
22
23
24
25
return $\mathcal{V}_{K}$, $\mathbf{m}_{K}$
Algorithm 2 Algorithm for learning $\mathbf{m}$ using Parity-aware Byte Pair Encoding with separate training and development sets.
Appendix B Intrinsic Tokenizer Evaluation Metrics
[…]
B.1 Vocabulary Usage
Vocabulary Utilization and Type-Token Ratio.
Vocabulary utilization measures the proportion of a tokenizer’s full vocabulary that is actively used when processing a given corpus. For tokenizer $T$ on corpus $\mathcal{D}$, we compute it as:
$\text{VocabUtil}(T)=\frac{|\{v:v\in\tau(\mathbf{b}),\mathbf{b}\in\mathcal{D}\}|}{|\mathcal{V}|}$ (8)
Here, the numerator counts the number of distinct tokens observed across the tokenization of all strings in the corpus. The type-token ratio quantifies lexical diversity by measuring the proportion of unique tokens (types) relative to the total number of tokens produced by a tokenizer:
$\text{TTR}(T)=\frac{|\{v:v\in\tau(\mathbf{b}),\mathbf{b}\in\mathcal{D}\}|}{\sum_{\mathbf{b}\in\mathcal{D}}|\tau(\mathbf{b})|}$ (9)
where $|\tau(\mathbf{b})|$ is the number of tokens produced by tokenizer $T$ for input $\mathbf{b}$. In words, the numerator counts distinct token types and the denominator counts total tokens across the corpus.
High vocabulary utilization and type-token ratio indicate efficient use of the learned vocabulary; low values of these metrics for a particular language may suggest tokenizer bias, as only a small portion of the tokenizer’s vocabulary is used/applicable for that language.
Average Token Rank.
Average token rank (Limisiewicz et al., 2023) measures the typical position of tokens in a tokenized text within the frequency-ordered vocabulary. In more detail, we compute the rank of each token (denoted as $\text{rank}(v)$) in our unigram frequency distribution $X_{T}$; rank 1 corresponds to the most frequent token. We compute average token rank across tokens in the evaluation corpus as:
$\text{AvgRank}(T)=\frac{\sum_{\mathbf{b}\in\mathcal{D}}\sum_{v\in\tau(\mathbf{b})}\text{rank}(v)}{\sum_{\mathbf{b}\in\mathcal{D}}|\tau(\mathbf{b})|}$ (10)
This metric can be seen as another measure of the proportion of the vocabulary used by a tokenizer. Lower average ranks indicate that the tokenizer predominantly uses a small set of tokens, while higher averages suggest more diverse token usage, including rare vocabulary items. When computed per language (i.e., when ranks are computed using the language’s respective frequency distribution), systematic differences in average token rank across languages reveal vocabulary allocation bias.
B.2 Information-theoretic Metrics
Compression Rate.
We evaluate compression rate—as defined in eq.˜2—across a parallel corpus. As discussed in §˜3.2, this enables us to use lines (documents) as our normalization unit. Recall that higher compression rates are generally desirable for computational efficiency in downstream tasks. In multilingual corpora, compression ratio disparities across languages indicate systematic tokenizer bias, where certain languages achieve better compression efficiency than others, potentially leading to unequal computational costs.
Rényi Entropy.
We compute Rényi entropy of order $\alpha$ over the empirical unigram frequency distribution $X_{T}$ for a given tokenizer $T$ to capture different aspects of token distribution:
$H_{\alpha}(X_{T})=\frac{1}{1-\alpha}\log_{2}\left(\sum_{v\in\mathcal{V}}p(v)^{\alpha}\right)$ (11)
for $\alpha\in\{1,2,\infty\}$. Rényi entropy provides a parametric family of measures that emphasize different aspects of the distribution: $H_{1}$ (Shannon entropy), $H_{2}$ (collision entropy), and $H_{\infty}$ (min-entropy). Rényi efficiency is Rényi entropy normalized by the size of the support, which is helpful for comparing tokenizers with different vocabulary sizes (Zouhar et al., 2023a). As all of our comparisons are between tokenizers of the same vocabulary size, we omit this normalization step and compare entropies directly.
B.3 Morphological and Multilingual Fairness Metrics
Fertility.
Fertility measures the average number of tokens produced per unit (word, character, or byte) by a tokenizer; the unit of interest for fertility is often the word, in which case, fertility quantifies how many tokens (on average) a word is broken up into. We use words as our normalization unit in our computations, as determined by the HuggingFace Whitespace Pretokenizer. We formally define tokenizer fertility for a given corpus $\mathcal{D}$ as:
$\text{Fertility}(T)=\frac{\sum_{\mathbf{b}\in\mathcal{D}}|\tau(\mathbf{b})|}{\sum_{\mathbf{b}\in\mathcal{D}}|\mathbf{b}|_{u}}$ (12)
This metric can give a sense for the computational efficiency imbued by a tokenizer, as well as for sequence length estimates for downstream modeling tasks.
MorphScore.
MorphScore (Arnett et al., 2025) evaluates tokenizer quality through morpheme-level precision and recall, measuring how well tokenizers preserve morphological information during segmentation. We point the reader to the origin Differences in cross-language MorphScore reveal how consistently a tokenizer’s sub-token boundaries align with true morpheme boundaries. A higher score in one language than another indicates that the tokenizer preserves that language’s morphological structure more faithfully. MorphScore provides a notion of both precision and recall (we point the reader to the original work for the exact description of the computation). Low precision indicates tokenizer oversegmentation; low recall is suggestive of under segmentation.
Tokenizer Fairness Gini Coefficient.
We adapt the Gini coefficient to measure fairness across languages by treating token costs (fertility values) as a distribution: Let $c_{1}\leq c_{2}\leq\ldots\leq c_{n}$ be the “costs” for languages $\mathcal{L}=\{l_{1},l_{2},\ldots,l_{n}\}$ incurred by a given tokenizer. Here, we quantify cost as the average number of tokens it takes to encode the unit of interest (e.g., a byte, word or line); when using a parallel corpus, this can be cost per line (document), which controls for discrepancies between average character byte length across different scripts. The Gini coefficient is then:
$\text{Gini}(T)=\frac{1}{n}\left(n+1-2\frac{\sum_{i=1}^{n}(n+1-i)c_{i}}{\sum_{i=1}^{n}c_{i}}\right)$ (13)
Values range from 0 (perfect equality) to 1 (maximum inequality). This metric condenses multilingual tokenizer fairness into a single number by measuring the degree of inequality in computational costs across languages; lower Gini coefficients indicate more equitable tokenizer compression across languages, while higher values suggest systematic bias toward certain languages.
Appendix C Additional Results and Ablation Studies
Figure 2: Compression rate of $128k$ tokenizers on the (unbalanced) 30-lang per language.
Figure 3: Vocabulary utilization of $128k$ tokenizers on the (unbalanced) 30-lang per language.
In this section, we present the results of our ablation studies. Table˜3 reports the intrinsic evaluation of tokenizers with a $128k$ vocabulary size on the (unbalanced) 60-lang dataset. Table˜4 shows the corresponding results for the (balanced) 30-lang dataset, also with a $128k$ vocabulary size. Finally, Table˜5 presents the intrinsic evaluation of tokenizers with a $256k$ vocabulary size on the (unbalanced) 30-lang dataset. Together, these results demonstrate the effectiveness of Parity-aware BPE across different language settings, vocabulary sizes, and data distributions.
Training Data Distribution.
To assess the sensitivity of the Parity-aware algorithm to training data distribution, we also analyze results for $128k$ tokenizers trained on the balanced version of the dataset. The results in Table˜4 indicate that parity-aware BPE tokenizers perform similarly to Classical BPE across most metrics. However, in terms of fertility, Classical BPE outperforms the parity-aware variants. This suggests that parity-aware tokenizers are particularly beneficial in unbalanced settings, where low-resource languages are more disadvantaged, whereas in balanced scenarios their advantage diminishes. We also interestingly see in Fig.˜4—again for the (balanced) 30-lang setting—that parity-aware tokenizers yield the largest absolute increases in vocabulary utilization for high-resource languages. Low- and medium-resource languages also improve, though to a smaller extent. One logical conclusion from this result is that the effect of parity-aware tokenizer is better described as balancing utilization across languages rather than directly compensating for data scarcity.
Script Analysis.
Fig.˜5 illustrates vocabulary utilization across languages for $128k$ tokenizers on the (unbalanced) 30-lang dataset. For Latin, Arabic, Hebrew, and Cyrillic scripts, the Parity-aware BPE (no-dev) variant outperforms all other parity-aware BPE versions, with Classical BPE ranking second. In contrast, for the CJK scripts, Classical BPE leads, while all parity-aware BPE variants perform similarly and closely follow. For the remaining scripts, the base Parity-aware BPE or the window-balanced variant significantly outperform other tokenizers. These findings suggest that different scripts benefit differently from parity-aware BPE approaches.
Tokenizer Type-Token Ratio Fertility Compression Rate Rényi Entropy ($\alpha$=2.5) Gini Coefficient MorphScore Precision MorphScore Recall Classical 0.0388 3.374 $\pm$ 0.027 0.0277 $\pm$ 0.0000 8.16 0.086 0.324 $\pm$ 0.031 0.407 $\pm$ 0.029 Parity-aware 0.0321 3.533 $\pm$ 0.029 0.0260 $\pm$ 0.0000 8.25 0.022 0.273 $\pm$ 0.030 0.379 $\pm$ 0.028 Parity-aware (hybrid) 0.0334 3.453 $\pm$ 0.028 0.0269 $\pm$ 0.0000 8.20 0.040 0.283 $\pm$ 0.030 0.379 $\pm$ 0.028 Parity-aware (window) 0.0392 3.438 $\pm$ 0.028 0.0270 $\pm$ 0.0000 8.17 0.030 0.305 $\pm$ 0.029 0.393 $\pm$ 0.028 Parity-aware (window+hybrid) 0.0409 3.362 $\pm$ 0.028 0.0278 $\pm$ 0.0000 8.12 0.044 0.317 $\pm$ 0.029 0.400 $\pm$ 0.028 Parity-aware (no-dev) 0.0401 3.412 $\pm$ 0.027 0.0276 $\pm$ 0.0000 8.16 0.080 0.334 $\pm$ 0.031 0.418 $\pm$ 0.029
Table 3: Intrinsic evaluation of $128k$ tokenizers on the (unbalanced) 60-lang dataset. Values are global statistics, except for MorphScore, which is macro-averaged across available languages.
Tokenizer Type-Token Ratio Fertility Compression Rate Rényi Entropy ($\alpha$=2.5) Gini Coefficient MorphScore Precision MorphScore Recall Classical 0.0780 4.175 $\pm$ 0.049 0.0307 $\pm$ 0.0001 8.09 0.050 0.409 $\pm$ 0.048 0.454 $\pm$ 0.046 Parity-aware 0.0765 4.207 $\pm$ 0.049 0.0300 $\pm$ 0.0001 8.12 0.011 0.405 $\pm$ 0.052 0.455 $\pm$ 0.049 Parity-aware (hybrid) 0.0767 4.192 $\pm$ 0.049 0.0303 $\pm$ 0.0001 8.11 0.016 0.404 $\pm$ 0.050 0.448 $\pm$ 0.048 Parity-aware (window) 0.0787 4.222 $\pm$ 0.050 0.0302 $\pm$ 0.0001 8.11 0.013 0.407 $\pm$ 0.050 0.455 $\pm$ 0.047 Parity-aware (window+hybrid) 0.0794 4.177 $\pm$ 0.049 0.0305 $\pm$ 0.0001 8.09 0.020 0.413 $\pm$ 0.049 0.457 $\pm$ 0.047 Parity-aware (no-dev) 0.0802 4.234 $\pm$ 0.050 0.0306 $\pm$ 0.0001 8.09 0.047 0.418 $\pm$ 0.048 0.463 $\pm$ 0.046 Parity-aware (hybrid+no-dev) 0.0800 4.231 $\pm$ 0.050 0.0307 $\pm$ 0.0001 8.08 0.048 0.415 $\pm$ 0.048 0.458 $\pm$ 0.046
Table 4: Intrinsic evaluation of $128k$ tokenizers on the (balanced) 30-lang dataset. Values are global statistics, except for MorphScore, which is macro-averaged across available languages.
Tokenizer Type-Token Ratio Fertility Compression Rate Rényi Entropy ($\alpha$=2.5) Gini Coefficient MorphScore Precision MorphScore Recall Classical 0.1239 3.767 $\pm$ 0.044 0.0340 $\pm$ 0.0001 7.85 0.052 0.515 $\pm$ 0.053 0.545 $\pm$ 0.051 Parity-aware 0.1205 3.809 $\pm$ 0.045 0.0334 $\pm$ 0.0001 7.87 0.010 0.498 $\pm$ 0.053 0.533 $\pm$ 0.051 Parity-aware (hybrid) 0.1212 3.803 $\pm$ 0.045 0.0336 $\pm$ 0.0001 7.86 0.012 0.506 $\pm$ 0.053 0.538 $\pm$ 0.051 Parity-aware (window) 0.1268 3.781 $\pm$ 0.045 0.0338 $\pm$ 0.0001 7.84 0.013 0.510 $\pm$ 0.052 0.543 $\pm$ 0.050 Parity-aware (window+hybrid) 0.1275 3.772 $\pm$ 0.045 0.0340 $\pm$ 0.0001 7.83 0.017 0.518 $\pm$ 0.052 0.548 $\pm$ 0.051 Parity-aware (no-dev) 0.1272 3.799 $\pm$ 0.044 0.0341 $\pm$ 0.0001 7.84 0.050 0.531 $\pm$ 0.052 0.559 $\pm$ 0.051 Parity-aware (hybrid+no-dev) 0.1271 3.797 $\pm$ 0.044 0.0341 $\pm$ 0.0001 7.84 0.050 0.531 $\pm$ 0.052 0.559 $\pm$ 0.051
Table 5: Intrinsic evaluation of $256k$ tokenizers on the (unbalanced) 30-lang dataset. Values are global statistics, except for MorphScore, which is macro-averaged across available languages.
Figure 4: Vocabulary utilization grouped by language resource levels for the $256k$ tokenizer trained on the (unbalanced) 30-lang dataset (left) and $128k$ tokenizer trained on the (balanced) 30-lang dataset (right).
Figure 5: Vocabulary utilization across language scripts for $128k$ tokenizers on the (unbalanced) 30-lang dataset.
Language Model Perplexities.
We report language model perplexities on the FineWeb2 validation set in Fig.˜6. Results are shown per language. We see a noticeably larger cross-lingual spread in perplexity for language models trained using the Classical BPE tokenizer than for those trained using Parity-aware variants. The Parity-aware tokenizers seem to eliminate the long tail present under Classical BPE while maintaining comparable mean perplexity across languages. Note that we normalize by number of bytes in the text rather than by number of tokens to account for differences in tokenization lengths.
Figure 6: Per-language perplexities normalized by byte of language models trained using the specified tokenizer trained on the (unbalanced) $128k$ 30-lang dataset. Results are computed on the language model validation set at the final checkpoint (see App.˜D for language model details).
Appendix D Language Model Training
Here we provide details about the language models used for evaluating extrinsic tokenizer metrics.
D.1 Model Architecture
We train models with 3 billion parameters ($3B$). All models follow the LLaMA architecture Touvron et al. (2023). The model size is determined by adjusting the number of layers, hidden sizes, and the number of attention heads.
D.2 Training Hyperparameters
We train our models using HuggingFace’s Nanotron trainer. Here we describe the hyperparameter selection for the different models’ training.
  • Learning Rate. We use a learning rate of 8e48e-4 with linear warmup on the first 4% of the training. Then we apply a “1-sqrt”-like cooldown for the last 20% of training  Hägele et al. (2024) as shown in Fig.˜7.
  • Optimizer. We use an AdamW Loshchilov and Hutter (2019) optimizer with $\beta=[0.9,0.95]$ for all our runs.
  • Weight Decay. We apply a weight decay $\lambda=0.1$ for regularization.
  • Batch Size. We fix our micro-batch size to 5 for all our runs.
02020404060608080100100022446688Tokens (B)Learning Rate ($\times 10^{-4}$)Learning Rate
Figure 7: Learning rate schedule over tokens with warmup and decay.
D.3 Hardware Setup
We train our models on a large-scale computing cluster consisting of nodes equipped with 4 NVIDIA Grace-Hopper H100 GPUs with 96 GB of memory. We train our 3B models on 64 nodes (or 256 GPUs) for around 18h per 100B tokens. Therefore our runs have a global batch size of 640 examples.
D.4 Sampling Methods
Let $\mathcal{L}$ be the set of languages in the dataset, and let $\pi^{\text{natural}}\in\Delta_{|\mathcal{L}|}$ represent the natural distribution of these languages, defined as:
$\pi_{l}^{\text{natural}}=\frac{\omega_{l}}{\sum_{l^{\prime}\in\mathcal{L}}\omega_{l^{\prime}}}$
where $\omega_{l}$ denotes the number of words (or tokens) for language $l$ in the dataset. In this work, we use the number of words as a proxy for language frequency, a common practice when presenting statistics for highly multilingual datasets Penedo et al. (2025). We use Temperature Sampling which is defined as:
This method adjusts the natural distribution using a temperature parameter $\tau$ to create a less skewed distribution:
$\pi_{l}^{\text{temp},\tau}=\frac{\omega_{l}^{1/\tau}}{\sum_{l^{\prime}\in\mathcal{L}}\omega_{l^{\prime}}^{1/\tau}}$
By tuning $\tau$, the distribution can be shifted towards uniformity, thereby reducing imbalance among languages.
Language Language Family Script Resource Level 30-lang 60-lang
English Indo-European (Germanic) Latin High
German Indo-European (Germanic) Latin High
French Indo-European (Romance) Latin High
Italian Indo-European (Romance) Latin High
Russian Indo-European (Slavic) Cyrillic High
Spanish Indo-European (Romance) Latin High
Japanese Japonic Kanji & Kana (CJK) Medium
Polish Indo-European (Slavic) Latin Medium
Portuguese Indo-European (Romance) Latin Medium
Vietnamese Austroasiatic Latin Medium
Turkish Turkic Latin Medium
Dutch Indo-European (Germanic) Latin High
Indonesian Austronesian Latin Medium
Arabic Afro-Asiatic (Semitic) Perso-Arabic Medium
Czech Indo-European (Slavic) Latin Medium
Persian (Farsi) Indo-European (Iranian) Perso-Arabic Medium
Greek Indo-European (Hellenic) Greek Medium
Chinese (Mandarin) Sino-Tibetan Hanzi (CJK) Medium
Hindi Indo-European (Indo-Aryan) Devanagari (Brahmic) Medium
Korean Koreanic Hangugeo (CJK) Medium
Thai Kra–Dai (Tai) Thai Medium
Hebrew Afro-Asiatic (Semitic) Hebrew Medium
Bengali Indo-European (Indo-Aryan) Bengali (Brahmic) Medium
Tamil Dravidian (Brahmic) Tamil Low
Georgian Kartvelian Georgian Low
Marathi Indo-European (Indo-Aryan) Devanagari (Brahmic) Medium
Filipino Austronesian Latin Low
Telugu Dravidian Telugu (Brahmic) Low
Norwegian Indo-European (Germanic) Latin Medium
North Azerbaijani Turkic Latin Low
Swedish Indo-European (Germanic) Latin Medium -
Romanian Indo-European (Romance) Latin Medium -
Ukrainian Indo-European (Slavic) Cyrillic Medium -
Hungarian Uralic (Ugric) Latin Medium -
Danish Indo-European (Germanic) Latin Medium -
Finnish Uralic (Finnic) Latin Medium -
Bulgarian Indo-European (Slavic) Cyrillic Low -
Slovak Indo-European (Slavic) Latin Low -
Catalan Indo-European (Romance) Latin Low -
Malay Austronesian Latin Low -
Urdu Indo-European (Indo-Aryan) Perso-Arabic Low -
Belarusian Indo-European (Slavic) Cyrillic Medium -
Basque Language Isolate Latin Low -
Tajik Indo-European (Iranian) Cyrillic Medium -
Sotho (Sesotho) Niger–Congo (Bantu) Latin Low -
Yoruba Niger–Congo Latin Low -
Swahili Niger-Congo (Bantu) Latin Low -
Estonian Uralic (Finnic) Latin Low -
Latvian Indo-European (Slavic) Latin Low -
Galician Indo-European (Romance) Latin Low -
Welsh Indo-European (Celtic) Latin Low -
Albanian Indo-European Latin Low -
Macedonian Indo-European (Slavic) Cyrillic Low -
Malayalam Dravidian Malayalam (Brahmic) Low -
Burmese Sino-Tibetan Mon–Burmese Low -
Gujarati Indo-European (Indo-Aryan) Gujarati (Brahmic) Low -
Afrikaans Indo-European (Germanic) Latin Low -
Hawaiian Austronesian Latin Low -
Northern Uzbek Turkic Latin Low -
Table 6: Details on the languages used to train and evaluate tokenizers.
Appendix E Downstream Benchmark Evaluation
We evaluate our models using HuggingFace’s Lighteval codebase Habib et al. (2023).
E.1 Benchmarks
We select 10 standard multilingual benchmarks to evaluate our models on various multilingual downstream tasks.
  • Belebele Bandarkar et al. (2024) – Multilingual reading comprehension dataset designed. It comprises passages and corresponding questions in multiple languages, aiming to assess the ability of models to comprehend and answer questions based on the provided texts.
  • mTruthfulQA Lin et al. (2022a); Lai et al. (2023) – Multilingual adaptation of the TruthfulQA benchmark. It consists of a wide range of questions aimed at detecting tendencies towards producing false or misleading information.
  • PAWS-X Yang et al. (2019) – Paraphrase identification and semantic similarity benchmark. It extends the original PAWS dataset to multiple languages, providing pairs of sentences with annotations indicating whether they are paraphrases.
  • XCodah Lin et al. (2021); Chen et al. (2019) – A dataset designed for evaluating adversarially-authored commonsense reasoning in natural language understanding. It extends the CODAH dataset to multiple languages.
  • XCSQA Lin et al. (2021); Talmor et al. (2019) Multilingual adaptation of the CommonsenseQA dataset, focusing on evaluating commonsense reasoning abilities across different languages. It consists of multiple-choice questions that require an understanding of common concepts and their relationships.
  • XNLI Conneau et al. (2018) – Designed to evaluate the ability of models to perform natural language inference (NLI) across multiple languages. It allows for the assessment of cross-lingual understanding and transfer learning capabilities in machine learning models.
  • XStoryCloze Mostafazadeh et al. (2017); Lin et al. (2022b) – Multilingual dataset to evaluate story comprehension and commonsense reasoning across different languages. It extends the StoryCloze Test to multiple languages, providing short stories with a missing ending and requiring models to choose the most appropriate conclusion from given options.
  • XWinogrande Sakaguchi et al. (2021); Muennighoff et al. (2023); Tikhonov and Ryabinin (2021) – Multilingual adaptation of WinoGrande, the adversarial version of the Winograd Schema Challenge. It consists of sentences with ambiguous pronouns that require models to correctly identify the antecedent based on contextual clues, assessing the model’s understanding of nuanced language and commonsense knowledge.
  • MMMLU Hendrycks et al. (2021); Lai et al. (2023) – Multilingual extension of MMLU, a benchmark designed to evaluate the performance of language models across a wide range of tasks.
  • INCLUDE Romanou et al. (2025) – Comprehensive knowledge- and reasoning-centric benchmark across 44 languages that evaluates multilingual LLMs for performance in the actual language environments where they would be deployed.
  • Exams Hardalov et al. (2020) – Dataset consisting of standardized test questions used to evaluate the problem-solving and reasoning abilities of language models. It includes questions from various subjects and educational levels, providing a measure of how well models can understand and generate responses to exam-style queries.
  • M3Exams Zhang et al. (2023) – Benchmark to evaluate the performance of language models on exam questions across different languages, subjects, and difficulty levels.
E.2 Score Aggregations
We aggregate benchmark results to compute a language-specific score for each model. Let $\mathcal{T}_{l}$ be the set of benchmarks (or tasks) containing a split for language $l$. The aggregated score for a model $m$ per language $l$ is defined as:
$s_{l}^{m}=\frac{1}{|\mathcal{T}_{l}|}\sum_{t\in\mathcal{T}_{l}}s_{t,l}^{m}$
where $s_{l}^{m}$ is the score of a model $m$ on the split $l$ of a task $t$ To mitigate biases arising from varying numbers of benchmarks per language, we compute a language-specific random baseline $\zeta_{l}$. This baseline helps assess whether a given aggregated score significantly outperforms random predictions. Specifically, we calculate the random baseline for each language as the average of the individual random baselines across all tasks that include language $l$:
$\zeta_{l}=\frac{1}{|\mathcal{T}_{l}|}\sum_{t\in\mathcal{T}_{l}}\zeta_{t}$
Language INCLUDE Belebele Exams M3Exam MMMLU mTruthfulQA PAWS-X XCodah XCOPA XCSQA XNLI XStoryCloze XWinoGrande
English - -
Chinese -
Vietnamese - - -
Arabic - - - -
German - - - - -
Spanish - - - -
French - - - -
Portuguese - - - - -
Hindi - - - - -
Russian - - -
Indonesian - - - - - - - -
Italian - - - -
Japanese - - - - - - - -
Swahili - - - - - -
Tamil - - - - - - - - - - -
Telugu - - - - - - - -
Thai - - - - - - - - -
Basque - - - - - - - - -
Turkish - - - - - - - -
Bulgarian - - - - - - - - -
Albanian - - - - - - - - - -
Polish - - - - - - - - - -
Bengali - - - - - - - - - -
Serbian - - - - - - - - - -
Estonian - - - - - - - - - - -
Macedonian - - - - - - - - - - -
Lithuanian - - - - - - - - - - -
Greek - - - - - - - - - - -
Urdu - - - - - - - - - - -
Catalan - - - - - - - - - - -
Persian - - - - - - - - - - - -
Finish - - - - - - - - - - - -
Korean - - - - - - - - - - - -
Quechua - - - - - - - - - - - -
Haitian Creole - - - - - - - - - - - -
Malay - - - - - - - - - - - -
Table 7: Coverage of downstream benchmarks across languages.